Circular No. 7179

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V382 VELORUM = NOVA VELORUM 1999 Additional photometry by A. C. Gilmore with the 0.6-m f/16 reflector

at Mount John University Observatory (cf. IAUC 7176; poor conditions with cirrus clouds): May 23.330 UT, V = 2.66(5), U − B = +0.06(2), B − V = +0.54(2), V − R = +0.34(2), V − I = +0.56(2) (airmass 1.015); May 24.312, 3.28(3), −0.38(1), +0.37(2), +0.38(3), +0.89(4).

Selected visual magnitude estimates: May 22.8958 UT, 3.0 (R. Levai, São Paulo, Brazil; independent discovery, forwarded by T. A. Napoleao); 23.917, 2.9 (T. A. Napoleao, São Paulo, Brazil); 24.334, 3.0 (D. A. J. Seargent, The Entrance, N.S.W.); 24.569, 3.4 (A. Pearce, Nedlands, W. Australia); 24.937, 3.2 (J. G. de S. Aguiar, Campinas, Brazil); 25.23, 3.8 (S. J. O’Meara, Volcano, HI); 25.43, 3.8 (R. H. McNaught, Bugaldie, N.S.W.).

COMET 95P/CHIRON D. C. Boice, National Astronomical Observatory of Japan, communi-

cates: “We have obtained millimeter-wave observations resulting in upper limits on the CO and HCN column abundances in the coma of (2060) = 95P/Chiron. These observations were made by T. Sekiguchi, N. Ukita, H. Hasegawa, and D. C. Boice during Feb. 10–18 UT using the 45-m radio telescope (+ AOS spectrometer) at Nobeyama Radio Observatory (NRO). Chiron was at mean distances of r = 9.32 AU and Δ = 9.40 AU during our observations. We observed the CO (J = 1–0) transition at 115.271204 GHz (2.6 mm) and the HCN (J = 1–0) transition at 88.6318473 GHz (3.4 mm). We achieved the following 3σ upper limits of (antenna temperature ∗ velocity), obtained from a total integration time of 7^{h}50^{m}: 0.043 K km/s for CO and 0.012 K km/s for HCN. Assuming that Chiron’s coma fills the NRO beam (18′′ for HCN and 14′′.3 for CO), and that the gas-excitation temperature is 8.8 K, these upper limits correspond to an average CO column density in the beam of < 2.1 × 10^{13} cm^{−2} and an average HCN column density of < 1.4 × 10^{10}. A Haser-model calculation, assuming an isotropic coma expansion velocity of 0.45 km/s and a CO linewidth twice the coma expansion velocity, implies upper limits of the gas-production rate of Q(CO) < 3.8 × 10^{28} s^{−1} and Q(HCN) < 3.3 × 10^{25} s^{−1}, if CO and HCN are parent molecules. If CO originates isotropically from subsurface layers, then the upper limit on the specific gas production rate is Z(CO) < 3.7 × 10^{13} cm^{−2} s^{−1}. Other searches for CO activity on Chiron are encouraged to aid in constraining models of activity on this peculiar object.”

1999 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green